For the fair 6-sided die, what is F(3)? F(7)? F(1.5)?
Answer #1

here as F(x) is cumulative distribution function which represent probability of getting x or lower then that

therefore F(x)=P(X<=x)

1)F(3)=P(X<=3)=P(X=1)+P(X=2)+P(X=3)=(1/6)+(1/6)+(1/6)=3/6 =1/2

2) F(7)=P(X<=7)=P(X=1)+P(X=2)+P(X=3)+P(X=4)+P(X=5)+P(X=6)+P(X=7)

=(1/6)+(1/6)+(1/6)+(1/6)+(1/6)+(1/6)+0 =1

3)F(1.5)=P(X<=1.5)=P(X=1)=1/6